OpenStudy (sleepyjess):

A ladder 5m long is leaning against a wall. If the lower end of the ladder slides away from the wall at the rate of .5 m/s, at what rate is the inclination theta of the ladder with respect to the ground changing when the lower end is 4m from the wall?

OpenStudy (sleepyjess):

@TheSmartOne @ganeshie8

OpenStudy (sleepyjess):



OpenStudy (sleepyjess):

Does this have to do with trig functions?

ganeshie8 (ganeshie8):

You may start with that diagram and use cosine ratio : \(\cos \theta = \dfrac{x}{5}\) where \(x\) is the bottom side

ganeshie8 (ganeshie8):

familiar with implicit differentiation ?

OpenStudy (sleepyjess):

Yes

ganeshie8 (ganeshie8):

Good. Implicitly differentiate above equation with respect to the time : \(t\) (Don't forget the chain rule)

ganeshie8 (ganeshie8):